A broker wants to sell a customer an investment costing $100 with an expected payoff in one year of $109.8. The customer indicates that a 9.8 percent return is not very attractive. The broker responds by suggesting the customer borrow $80 for one year at 7.8 percent interest to help pay for the investment.

a. What is the customer’s expected return if she borrows the money? (Round your answer to 1 decimal place.)
Answer #1

The expected return is calculated below:

100 – 80 = 109.8 1+i 80 x 1.078 1+i

109.8-86.24 20 = 1+1

1+1= 23.56 20

i = 0.178

Expected return is 17.80%.